# Synthesis and stability of biomolecules under Earth’s upper mantle conditions

### v1

@@ III.1 First Principles molecular dynamics simulations (FPMD)
Free-energy calculations by coupling FPMD simulations with enhanced sampling
@@
To compute the free-energy landscape for compound formation, we employ the adaptive biasing force method (ABF) darve2008adaptive as implemented in the software package SSAGES sidky2018ssages coupled with the FPMD software Qbox. Different from other advanced sampling methods that add biases to the energy of the system, this method seeks a flattening of the generalized force. The basic ABF algorithm is shown as follows, the running average of the force in bin k along the reaction coordinate space $\xi$ could be calculated:
${{F}_{\xi}}\left({{N}_{step,}}k\right)=\frac{1}{\nu\left({{N}_{step,}}k\right)}\sum\limits_{i=1}^{\nu\left({{N}_{step,}}k\right)}{{{F}_{i}}}\left(t_{i}^{k}\right)$
${{F}_{i}}(t_{i}^{k})=\frac{d}{dt}\left({{M}_{\xi}}\frac{d\xi}{dt}\right)\left|{}_{t_{i}^{k}}\right.$
@@
[…]
In free-energy calculations, similar to the FPMD simulations described above, we utilize enhanced sampling simulations in the NVT ensemble, employing the BDP thermostat with a time step of 0.24 fs, using the PBE exchange-correlation functional with plane waves (a 65 Ry energy cutoff), and the ONCV pseudopotential. For each specific calculation, we use different numbers of grids for distance variables to store local estimates of the mean derivative of the free energy surface. Furthermore, we select different kinds of distance between atoms as collective variables (CVs) for different organic compounds.
@@
Details of simulations, percentage of the C-N containing species with time, hydrocarbon products and free energy calculation, another strategies of formation of ribose and uracil, and their rings.
@@ Acknowledgements
The authors would like to acknowledge support from the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16307618, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China (Project 11774072 and Excellent Young Scientists Fund), the Alfred P. Sloan Foundation through the Deep Carbon Observatory, and the Hetao Shenzhen/Hong Kong Innovation and Technology Cooperation (HZQB-KCZYB-2020083).

### v2

@@ III.1 Ab initio molecular dynamics simulations (AIMD)
Free-energy calculations by coupling AIMD simulations with enhanced sampling
@@
To compute the free-energy landscape, we employed the adaptive biasing force method (ABF) darve2008adaptive as implemented in the software package SSAGES sidky2018ssages coupled with Qbox code. Unlike other advanced sampling methods that introduce biases into the energy, this method aims to flatten the generalized force. The running average of the force in the $k$th bin along the reaction coordinate $\xi$ is calculated as:
$F_{\xi}(N_{\text{step}},k)$ $=\frac{1}{\nu(N_{\text{step}},k)}\sum_{i=1}^{\nu(N_{\text{step}},k)}F_{i}(t_{i}^{k})$ (1)
$F_{i}(t_{i}^{k})$ $=\frac{d}{dt}\left(M_{\xi}\frac{d\xi}{dt}\right)\bigg{|}_{t_{i}^{k}}$ (2)
where $\nu\left({{N}_{step}},k\right)$ is the number of samples collected in the $k$th bin after $N_{step}$ steps, ${{F}_{i}}\left(t_{i}^{k}\right)$ is the $i$th force sample at the time $t_{i}^{k}$, M is the mass matrix. The bias force -$\text{-}\left(\nabla\xi\right){{F}_{\xi}}\left({{N}_{step,}}k\right)$ is added to the calculation until the free energy landscape is flat. We finally integrated the total biased force to get the free energy. For example, the free energy $\Delta{{A}_{a\to b}}$ between two states $a$ and $b$ is calculated as:
@@
Details of the simulations, the percentage of C-N containing species over time, the hydrocarbon products, the free energy calculations, other strategies for the formation of ribose and uracil, and their ring structures, compositions of fluids at different oxygen fugacity.
@@ Acknowledgements
This work was supported by the Croucher Foundation through the Croucher Innovation Award, Hong Kong Research Grants Council (Projects GRF-16301723, GRF-16306621, and C6021-19EF), National Natural Science Foundation of China through the Excellent Young Scientists Fund (22022310), and the Hetao Shenzhen/Hong Kong Innovation and Technology Cooperation (HZQB-KCZYB-2020083). R.T. would like to acknowledge support from the Original Exploration Plan-Key project (42150104). Contributions by D.A.S. constitute material based upon work supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Geosciences program under Award Number DE-SC0019830 as well as NSF Petrology and Geochemistry Grant Number 2032039. Part of this work was carried out using computational resources from the National Supercomputer Center in Guangzhou, China.